# Score Distillation Sampling with Learned Manifold Corrective

### v1

@@ Abstract
Score Distillation Sampling (SDS) is a recent but already widely popular method that relies on an image diffusion model to control optimization problems using text prompts. In this paper, we conduct an in-depth analysis of the SDS loss function, identify an inherent problem with its formulation, and propose a surprisingly easy but effective fix. Specifically, we decompose the loss into different factors and isolate the component responsible for noisy gradients. In the original formulation, high text guidance is used to account for the noise, leading to unwanted side effects. Instead, we train a shallow network mimicking the timestep-dependent denoising deficiency of the image diffusion model in order to effectively factor it out. We demonstrate the versatility and the effectiveness of our novel loss formulation through several qualitative and quantitative experiments, including optimization-based image synthesis and editing, zero-shot image translation network training, and text-to-3D synthesis.
@@ 1 Introduction
Image diffusion models [11] have recently become the de facto standard for image generation. Especially text-to-image models [26, 33, 30, 29] have emerged as powerful tools for high fidelity, diverse, image synthesis. By being controlled through natural language, these models are extremely easy to use and thus open up a wide range of creative applications, without the need for special training. Beyond image synthesis, image diffusion models have been successfully deployed in applications like image restoration, inpatining, editing, super-resolution, or colorization [4, 13, 24, 7], among others. Image diffusion models are typically obtained using very large training sets and thus inherently represent the distribution of natural images. While applications typically make use of the generation process of diffusion models, e.g. by inserting images into the process [24, 23] or by altering the denoising function [5, 4, 13], relatively little research has been conducted on how diffusion models can be uses as rich, general purpose image priors. Score Distillation Sampling (SDS) proposed in DreamFusion [28] is one of the few exceptions. The SDS loss formulation uses a pretrained text-to-image model [33] to measure how well a given text matches an observation. In DreamFusion this loss is being used to generate 3D assets from textual descriptions, an idea that was quickly adopted [36, 19, 31, 25, 22, 16] and essentially established a new research direction: text-to-3D synthesis. While being mostly used in this context, the design of the SDS loss is by no means only tailored to text-to-3D applications. In fact, SDS is an image loss and can be used in much wider contexts [12, 8]. However, in its original formulation the SDS loss may degrade the observation, be too eager to match the text prompt, or provide meaningless gradients which inject noise into the objective. In this paper, we conduct an extensive analysis of the SDS loss, identify an inherent problem with its formulation, propose a surprisingly easy but effective fix, and demonstrate its effectiveness in several applications, including optimization-based image synthesis and editing, zero-shot training of image translation networks, and text-to-3D synthesis. Concretely, our new loss formulation – Score Distillation Sampling with Learned Manifold Corrective, or LMC-SDS for short – aims to provide better gradients along the direction of the learned manifold of real images. We present evidence that gradients towards the learned manifold are extremely noisy in SDS and that high text guidance is needed to compensate for the noisy signal. Further, we show that high guidance – or lack thereof – is a possible explanation for the artifacts observed with SDS, cf. Sec. 3. In contrast, applications relying on our novel formulation benefit from meaningful manifold gradients, may use lower text guidance, and produce results of overall higher visual fidelity.
@@ 2 Related Work
The first class of methods leverage image diffusion networks to solve specific tasks by relying on the diffusion sampling process. We discuss here a number of representative methods and refer the reader to [18, 6] for detailed discussions. DDRM [13] tackles linear inverse problems such as deblurring, super-resolution, colorization, or inpainting by defining a modified variational posterior distribution conditioned on the observation signal, which can then be used for sampling. DPS [4] focuses on general inverse problems and proposes to augment the score prediction at each denoising step with an additional term that models the likelihood of the observations. MCG [5] builds on top of [4] and improves the fidelity of generation by introducing additional manifold constraints. SDEdit [24] uses diffusion models for image editing by first adding noise to the input image and then denoising it using the text-to-image diffusion prior. RePaint [23] uses unconditional image diffusion models as a prior for solving image inpainting. All these methods use pre-trained diffusion models. In contrast, methods like Palette [32] or Imagic [14] require training a diffusion model from scratch or fine-tuning a pre-trained one, respectively. This line of work exclusively deals with image editing problems and is typically restricted to the resolution of the diffusion models.
@@
More related to our method is the second category: In an orthogonal strategy this body of work uses pre-trained diffusion models within general purpose loss functions in iterative optimization settings. Score Jacobian Chaining [36] propagates the score of a text-to-image diffusion model through a differentiable renderer to supervise the generation of 3D models. DreamFusion [28] takes an orthogonal approach and proposes the Score Distillation Sampling loss. SDS is a modification of the diffusion training objective that does not require expensive backpropagation through the diffusion model and encourages the alignment of the generated image with a conditioning signal. DDS [8] proposes an improved version of SDS, specifically designed for image editing and aims to reduce the artifacts introduced by noisy SDS gradients. In contrast to DDS, our method does not require a source text prompt. SparseFusion [39] introduces a multi-step SDS loss that performs multiple denoising steps instead of one, which adds significant computational overhead. ProlificDreamer [37] proposes a generalization of SDS from point estimates to distributions, but unlike our approach requires computationally expensive fine-tuning of a diffusion model during optimization. Collaborative Score Distillation [15] generalizes SDS to multiple samples and is inspired by Stein variational gradient descent [21]. Category Score Distillation Sampling (C-SDS) focuses on 3D generation and replaces the noise estimation error in SDS with a difference between the noise predictions of a standard and a multi-view consistent diffusion model [40]. Unlike our method, C-SDS is only applicable to 3D generation tasks. In contrast, we propose a general purpose loss function with very little computational overhead.
@@ 3 Analysis
where $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ refers to the noisy version of $\mathbf{z}$ and $w(t)$ is a weighting function, omitted in the sequel. DreamFusion [28] showed that given a pre-trained model $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}$ the diffusion loss can be utilized for optimization problems. For an arbitrary differentiable rendering function returning $\mathbf{z}$ and parameterized by $\boldsymbol{\theta}$, the gradient of the denoising function w.r.t. $\boldsymbol{\theta}$ is given by:
$\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{diff}}=\left(\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\phi}}(\mathbf{z}_{t},\boldsymbol{y},t)-\boldsymbol{\epsilon}\right)\frac{\partial\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\phi}}(\mathbf{z},\boldsymbol{y},t)}{\partial\mathbf{z}_{t}}\frac{\partial\mathbf{z}_{t}}{\partial\boldsymbol{\theta}}$ (2)
@@
$\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\phi}}(\mathbf{z},\boldsymbol{y},t)=\omega\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},\boldsymbol{y},t)+(1-\omega)\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},t),$ (4)
which we can rewrite as:
$\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\phi}}(\mathbf{z},\boldsymbol{y},t)=\omega\left(\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},\boldsymbol{y},t)-\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},t)\right)+\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},t).$ (5)
By inserting Eq. 5 into Eq. 3 we can express $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{SDS}}$ as:
@@
$\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{cond}}=\omega\left(\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},\boldsymbol{y},t)-\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},t)\right)\frac{\partial\mathbf{z}_{t}}{\partial\boldsymbol{\theta}}$ (6b)
$\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}=\left(\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},t)-\boldsymbol{\epsilon}\right)\frac{\partial\mathbf{z}_{t}}{\partial\boldsymbol{\theta}}.$ (6c)
@@
DreamFusion [28] originally proposed to set $\omega=100$. This means that, following our formulation, $\mathcal{L}_{\text{cond}}$ is weighted proportionally much higher than $\mathcal{L}_{\text{proj}}$. This can be problematic as large guidance weights have been identified to be an important cause of over-exposure in diffusion models [20]. In fact, DreamFusion produces detailed results but also has a tendency towards unrealistically saturated colors, see also Fig. 9. On the other hand, setting $\omega$ to lower values has been reported to result in very blurry images [8], a behavior that we also demonstrate in Sec. 5. Both behaviors, (1) the tendency to produce overly saturated results when using high guidance as well as (2) blurry results for low guidance, can be explained by looking at the components of $\mathcal{L}_{\text{SDS}}$ in isolation. To this end, we consider the case where the rendering function is a no-op and simply returns pixel values, i.e. $\mathbf{z}=\boldsymbol{\theta}$. In Fig. 1, we visualize $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{cond}}$ and $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}$ for this case. $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{cond}}$ returns the direction towards an image with altered colors. However, as $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{cond}}$ is not anchored in the manifold of realistic images, the gradient may eventually point away from that manifold. $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{cond}}$ will always point towards changes that “better” respect the prompt and eventually produce saturated colors or other artifacts (cf. Sec. 5). $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}$, on the other hand, incorrectly marks high-frequent detail originating from $\mathbf{z}$ for removal – a reason for blurriness.
@@
We conclude the following: (1) The reason for oversaturated colors and other artifacts is due to $\mathcal{L}_{\text{cond}}$ being “over-eager” to maximize the agreement of the image with the conditioning variable. (2) $\mathcal{L}_{\text{proj}}$ provides deficient gradients ultimately removing high-frequent detail instead of “correcting” the gradient towards on-manifold solutions. $\mathcal{L}_{\text{proj}}$ appears to be a non-optimal choice and we provide an alternative in the sequel. But first we use our decomposition to discuss other approaches to “fix” the SDS loss.
@@
DDS [8] proposes to use the difference between two SDS losses for image editing, one computed on the current state w.r.t. the target prompt and one computed on the initial image w.r.t. a textual description of that image. By computing an SDS difference, DDS entirely removes $\mathcal{L}_{\text{proj}}$ and introduces a negative gradient pointing away from the initial state. While the root for blurry results is removed by eliminating $\mathcal{L}_{\text{proj}}$, the optimization is no longer anchored along the learned manifold. Additionally, an initial image is needed, limiting possible use-cases. In a different approach, other methods aim for a better noise prediction, where by improving $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},t)$, one may also improve the underlying $\mathcal{L}_{\text{proj}}$, as the difference between the predicted and true noise. E.g. SparseFusion [39] introduces multi-step SDS. Hereby, instead of denoising in one step, up to 50 steps are used. While this may produce more informative gradients in $\mathcal{L}_{\text{proj}}$, this is also significantly more expensive. Through different means but towards the goal of better noise prediction, concurrent work [37] learns a low-rank adaptation of the text-to-image model. While this approach is effective, it also adds the overhead of fine-tuning an additional diffusion model. In a third vein, some recent methods aim to alleviate artifacts introduced by SDS by advanced schedules for sampling $t$ or by introducing negative prompts [1, 34]. In contrast, we aim for a generally applicable solution.

### v2

@@ Abstract
Score Distillation Sampling (SDS) is a recent but already widely popular method that relies on an image diffusion model to control optimization problems using text prompts. In this paper, we conduct an in-depth analysis of the SDS loss function, identify an inherent problem with its formulation, and propose a surprisingly easy but effective fix. Specifically, we decompose the loss into different factors and isolate the component responsible for noisy gradients. In the original formulation, high text guidance is used to account for the noise, leading to unwanted side effects such as oversaturation or repeated detail. Instead, we train a shallow network mimicking the timestep-dependent frequency bias of the image diffusion model in order to effectively factor it out. We demonstrate the versatility and the effectiveness of our novel loss formulation through qualitative and quantitative experiments, including optimization-based image synthesis and editing, zero-shot image translation network training, and text-to-3D synthesis.
@@ 1 Introduction
Image diffusion models [10] have recently become the de facto standard for image generation. Especially text-to-image models [27, 34, 31, 30] have emerged as powerful tools for high fidelity, diverse, image synthesis. By being controlled through natural language, these models are extremely easy to use and thus open up a wide range of creative applications, without the need for special training. Beyond image synthesis, image diffusion models have been successfully deployed in applications like image restoration, inpatining, editing, super-resolution, or colorization [4, 14, 25, 7], among others. Image diffusion models are typically obtained using very large training sets and thus inherently represent the distribution of natural images. While applications typically make use of the generation process of diffusion models, e.g. by inserting images into the process [25, 24] or by altering the denoising function [5, 4, 14], relatively little research has been conducted on how diffusion models can be used as rich, general purpose image priors. Score Distillation Sampling (SDS) proposed in DreamFusion [29] is one of the few exceptions. The SDS loss formulation uses a pretrained text-to-image model [34] to measure how well a given text matches an observation. In DreamFusion this loss is being used to generate 3D assets from textual descriptions, an idea that was quickly adopted [37, 20, 32, 26, 23, 17] and essentially established a new research direction: text-to-3D synthesis. While being mostly used in this context, the design of the SDS loss is by no means only tailored to text-to-3D applications. In fact, SDS is an image loss and can be used in much wider contexts [12, 8]. However, in its original formulation the SDS loss may degrade the observation, be too eager to match the text prompt, or provide meaningless gradients which inject noise into the objective. In this paper, we conduct an extensive analysis of the SDS loss, identify an inherent problem with its formulation, propose a surprisingly easy but effective fix, and demonstrate its effectiveness in several applications, including optimization-based image synthesis and editing, zero-shot training of image translation networks, and text-to-3D synthesis. Concretely, our new loss formulation – Score Distillation Sampling with Learned Manifold Corrective, or LMC-SDS for short – aims to provide better gradients along the direction of the learned manifold of real images. We present evidence that gradients towards the learned manifold are extremely noisy in SDS and that high text guidance is needed to compensate for the noisy signal. Further, we show that high guidance – or lack thereof – is a possible explanation for the artifacts observed with SDS, cf. Sec. 3. In contrast, applications relying on our novel formulation benefit from meaningful manifold gradients, may use lower text guidance, and produce results of overall higher visual fidelity.
@@ 2 Related Work
The first class of methods leverage image diffusion networks to solve specific tasks by relying on the diffusion sampling process. We discuss here a number of representative methods and refer the reader to [19, 6] for detailed discussions. DDRM [14] tackles linear inverse problems such as deblurring, super-resolution, colorization, or inpainting by defining a modified variational posterior distribution conditioned on the observation signal, which can then be used for sampling. DPS [4] focuses on general inverse problems and proposes to augment the score prediction at each denoising step with an additional term that models the likelihood of the observations. MCG [5] builds on top and improves the fidelity of generation by introducing additional manifold constraints. SDEdit [25] uses diffusion models for image editing by first adding noise to the input image and then denoising it using the text-to-image diffusion prior. RePaint [24] uses unconditional image diffusion models as a prior for solving image inpainting. All these methods use pre-trained diffusion models. In contrast, methods like Palette [33] or Imagic [15] require training a diffusion model from scratch or fine-tuning a pre-trained one, respectively. This line of work exclusively deals with image editing problems and is typically restricted to the resolution of the diffusion models.
@@
More related to our method is the second category: This body of work uses pre-trained diffusion models within general purpose loss functions in iterative optimization settings. Score Jacobian Chaining [37] propagates the score of a text-to-image diffusion model through a differentiable renderer to supervise the generation of 3D models. DreamFusion [29] takes an orthogonal approach and proposes the Score Distillation Sampling loss. SDS is a modification of the diffusion training objective that does not require expensive backpropagation through the diffusion model and encourages the alignment of the generated image with a conditioning signal. DDS [8] proposes an improved version of SDS, specifically designed for image editing and aims to reduce the artifacts introduced by noisy SDS gradients. In similar spirit and in concurent work, NFSD [13] aims to reduce noise in the gradients by adding negative conditioning. In contrast, our method neither requires a source prompt nor a negative prompt. SparseFusion [40] introduces a multi-step SDS loss that performs multiple denoising steps instead of one, which adds significant computational overhead. ProlificDreamer [38] proposes a generalization of SDS from point estimates to distributions, but unlike our approach requires computationally expensive fine-tuning of a diffusion model during optimization. Collaborative Score Distillation [16] generalizes SDS to multiple samples and is inspired by Stein variational gradient descent [22]. Category Score Distillation Sampling (C-SDS) focuses on 3D generation and replaces the noise estimation error in SDS with a difference between the noise predictions of a standard and a multi-view consistent diffusion model [41]. Unlike our method, C-SDS is only applicable to 3D generation tasks. In contrast, we propose a general purpose loss function with very little computational overhead.
@@ 3 Analysis
where $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ refers to the noisy version of $\mathbf{z}$ and $w(t)$ is a weighting function, omitted in the sequel. DreamFusion [29] showed that given a pre-trained model $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}$ the diffusion loss can be utilized for optimization problems. For an arbitrary differentiable rendering function returning $\mathbf{z}$ and parameterized by $\boldsymbol{\theta}$, the gradient of the denoising function w.r.t. $\boldsymbol{\theta}$ is given by
$\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{diff}}=\left(\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\phi}}(\mathbf{z}_{t},\boldsymbol{y},t)-\boldsymbol{\epsilon}\right)\frac{\partial\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\phi}}(\mathbf{z},\boldsymbol{y},t)}{\partial\mathbf{z}_{t}}\frac{\partial\mathbf{z}_{t}}{\partial\boldsymbol{\theta}}.$ (2)
@@
$\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\phi}}(\mathbf{z}_{t},\boldsymbol{y},t)=\omega\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},\boldsymbol{y},t)+(1-\omega)\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t),$ (4)
which we can rewrite as
$\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\phi}}(\mathbf{z}_{t},\boldsymbol{y},t)=\omega\left(\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},\boldsymbol{y},t)-\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t)\right)+\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t).$ (5)
By inserting Eq. 5 into Eq. 3 we can express $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{SDS}}$ as
@@
$\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{cond}}=\omega\left(\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},\boldsymbol{y},t)-\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t)\right)\frac{\partial\mathbf{z}_{t}}{\partial\boldsymbol{\theta}}$ (6b)
$\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}=\left(\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t)-\boldsymbol{\epsilon}\right)\frac{\partial\mathbf{z}_{t}}{\partial\boldsymbol{\theta}}.$ (6c)
@@
DreamFusion [29] originally proposed to set $\omega=100$. This means that, following our derivation, $\mathcal{L}_{\text{cond}}$ is weighted proportionally much higher than $\mathcal{L}_{\text{proj}}$. This can be problematic as large guidance weights have been identified to be an important cause of over-exposure in diffusion models [21]. In fact, DreamFusion produces detailed results but also has a tendency towards unrealistically saturated colors, see also Fig. 8. On the other hand, setting $\omega$ to lower values has been reported to result in very blurry images [8], a behavior that we also demonstrate in Sec. 5. Both behaviors, (1) the tendency to produce overly saturated results when using high guidance as well as (2) blurry results for low guidance, can be explained by looking at the components of $\mathcal{L}_{\text{SDS}}$ in isolation. To this end, we consider the case where the rendering function is a no-op and simply returns pixel values, i.e. $\mathbf{z}=\boldsymbol{\theta}$. In Fig. 1 (left), we visualize $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{cond}}$ and $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}$ for this case. $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{cond}}$ returns the direction towards an image with altered colors. However, as $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{cond}}$ is not anchored in the manifold of natural images, the gradient may eventually point away from that manifold. $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{cond}}$ will always point towards changes that “better” respect the prompt and eventually produce saturated colors or other artifacts (cf. Sec. 5). $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}$, on the other hand, incorrectly marks high-frequency detail originating from $\mathbf{z}$ for removal – the reason for blurriness. This behavior can be understood when rewriting $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}$ as
$\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}=\left(\frac{\sqrt{\alpha_{t}}}{\sqrt{1-\alpha_{t}}}\left(\mathbf{z}-\mathbf{z}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t)\right)\right)\frac{\partial\mathbf{z}_{t}}{\partial\boldsymbol{\theta}},$ (7)
which is equivalent to comparing the predicted and true noise. By looking at Eq. 7, we can now see a possible explanation why $\mathcal{L}_{\text{proj}}$ might not be optimal: $\hat{\mathbf{z}}_{t}=\mathbf{z}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t)$ is an approximation of $\mathbf{z}$ and most probably not a perfect replica of $\mathbf{z}$, even when $\mathbf{z}$ already lies on the natural image manifold, see Fig. 1 (left). This is especially the case for large values of $t$, where the denoising model has to reconstruct the true image from almost pure noise. More formally, we can observe a frequency bias in $\mathbf{z}_{\boldsymbol{\phi}}$ dependent on $t$ when comparing images $\mathbf{z}$ with their denoised counterparts $\hat{\mathbf{z}}_{t}$, see Fig. 1 (right).
@@
We conclude the following: (1) The reason for oversaturated colors and other artifacts is due to $\mathcal{L}_{\text{cond}}$ being “over-eager” to maximize the agreement of the image with the conditioning variable. (2) $\mathcal{L}_{\text{proj}}$ provides deficient gradients ultimately removing high-frequency detail instead of “correcting” the gradient towards on-manifold solutions due to a frequency bias in $\mathbf{z}_{\boldsymbol{\phi}}$. Thus, $\mathcal{L}_{\text{proj}}$ appears to be a non-optimal choice and we provide an alternative in the sequel. But first we use our decomposition to discuss other approaches to “fix” the SDS loss.
@@
DDS [8] proposes to use the difference between two SDS losses for image editing, one computed on the current state w.r.t. the target prompt and one computed on the initial image w.r.t. a textual description of that image. By computing an SDS difference, DDS entirely removes $\mathcal{L}_{\text{proj}}$ and introduces a negative gradient pointing away from the initial state. However, the need for an initial image is limiting possible use-cases. In concurrent work, NFSD [13] aims to obtain better results by replacing $\mathcal{L}_{\text{proj}}$ with a $\mathcal{L}_{\text{cond}}$ variant using negative prompts for most timesteps, and a loss based on $\boldsymbol{\epsilon}$ for small $t$. While the root for blurry results is removed by eliminating $\mathcal{L}_{\text{proj}}$, both DDS and NFSD do not anchor the optimization along the learned manifold. Other methods aim for a better noise prediction, where by improving $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},t)$, one may also improve $\mathcal{L}_{\text{proj}}$ ($\mathcal{L}_{\text{proj}}$ is the difference between the predicted and true noise). E.g. SparseFusion [40] introduces multi-step SDS. Hereby, instead of denoising in one step, up to 50 steps are used. While this may produce more informative gradients in $\mathcal{L}_{\text{proj}}$, this is also significantly more expensive. Through different means but towards the same goal, the concurrent work ProlificDreamer [38] learns a low-rank adaptation of the text-to-image model. While this approach is effective, it also adds the overhead of fine-tuning an additional diffusion model. Finally, some recent methods aim to alleviate artifacts by augmenting the original SDS formulation with advanced schedules for sampling $t$ or with negative prompts [1, 35]. In contrast, we aim for a generally applicable solution.